A rectangle has an area of 432 square inches. The width of the rectangle is 3/4 times its length. Find the dimensions of the rectangle
Answer by mananth(16946)

A rectangle has an area of 432 square inches. The width of the rectangle is 3/4 times its length. Find the dimensions of the rectangle
length =x
width = 3x/4
3x/4 * x= 432
3x^2/4=432
3x^2=1728
x^2=1728/3
x^2=576
x=24 in the length
width = 3/4 *24
=18 in
...
m.ananth@hotmail.ca